OpenStudy (anonymous):

how do you simplify: -6sqrt(9x)+3sqrt(64x)-sqrt(50x)

OpenStudy (anonymous):

You can break up the square roots to \[-6\sqrt{x}\sqrt{9} + 3\sqrt{x}\sqrt{64} -\sqrt{x}\sqrt{50}\] sqrt(9) = 3 sqrt(64) = 6 sqrt(50) = 5sqrt(2) So now we have: \[-18\sqrt{x} + 18\sqrt{x} - 5\sqrt{2x}\] This simplifies to -5sqrt(2x)

OpenStudy (anonymous):

but...i thought the square root of 64 was 8?

OpenStudy (anonymous):

whoops, typo. It is.

OpenStudy (anonymous):

Instead \[-18\sqrt{x} + 24\sqrt{x} - 5\sqrt{2x}\] so it simplifies to: 6sqrt(x) - 5sqrt(2x)

OpenStudy (anonymous):

Thank you so much!